Log 2=0.30 log3=0.477 log5^1/3and also the value of log 108

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
log 5^1/3 = 1/3 log 5 = 1/3 log (10/2) = 1/3 ( log 10 – log 2) = 1/3 ( 1 – 0.3010) = 0.699/3 = 0.233
 
Hope it helps
 
Regards
Arun